12. (5 points) A What is the magnitude of 2.20 A is flowing through a 1000-turn solenoid of length 35.0 cm. i.e. the turns perpendicular field inside the solenoid? If this solenoid perfect one, 5 degree, what with axis the (90 degree), the angle from turns to axis is is the magnitude of the magnetic field inside of the solenoid?
Answer #1

for ideal solenoid

B= \mu ni

where n is number of coil per unit length

B= 4*3.14*10^-7 * (1000/.35)*2.20

=7.89 \mu T ans

IN Case of imperfect solenoid

B= \mu ni cos\theta= 7.89 * cos (5) = 7.85 \mu T   ans